# Gradient-free neural topology optimization

### v1

@@ 3.2 Benchmarking set-up
We develop and benchmark our method considering a standard topology optimization problem, namely compliance minimization with a volume constraint. While this problem can be easily solved using gradient-based approaches, it is not the ultimate application of the proposed method. Nevertheless, it provides a good benchmark for the development of the method because, first, the "ground truth" solutions can be easily obtained using a gradient-based approach, allowing us to verify ’how far’ our method is from a good solution. Secondly, these problems are relatively cheap to evaluate, as opposed to design problems based on more complex physics.
@@
To provide statistically meaningful performance measures, we test our optimization framework on two sets of optimization problems. The first set contains 25 in-distribution design problems (sets of boundary conditions). These problems were generated in the same way as the problems seen during the training (i.e. with random point loads and point supports). The second set contains 25 out-of-distribution problems, generated using the approach of Hoyer et al. [12] (the same as the testing dataset). These problems include distributed loads and supports, therefore they provide a better insight into the generalization of the model.
@@
To quantify the performance of different model and optimizer configurations across the whole problem set, for each problem, we normalize the compliance value with the compliance of the MMA solution (considered as the ’ground truth’), and in this way calculate the relative error. Furthermore, to account for stochastic factors within the optimization process (e.g. initialization of the optimizer), the optimization process for each configuration and each problem is repeated with 15 different random seeds. The performance is evaluated for grayscale and binary (thresholded) designs. For thresholding we apply the simplest strategy from [50] in a post-processing step. In that case, the relative error is calculated with respect to the value obtained for the thresholded MMA design.
@@
For gradient-free optimizer, we use Covariance Matrix Adaptation Evolutionary Strategies (CMA-ES) [44, 51] and its Bi-Population variant (BIPOP-CMA-ES) [52]. The CMA-ES and its variants (such as BIPOP-CMA-ES) are currently regarded as state-of-the-art gradient-free optimization algorithms [7]. Additionally, CMA-ES was considered in a similar setting [53]. Notwithstanding, we considered several alternative gradient-free optimizers to ensure that we selected the most favorable one for our problems, namely Particle Swarm Optimization [54], Differential Evolution [55], Evolutionary Strategies meta-optimizer [56], and Simple Genetic Algorithm [57]. However, even with hyperparameter tuning these optimizers exhibited considerably worse performance than CMA-ES within the assigned budget.
@@
The volume constraint is enforced using a constrained sigmoid transformation, following the approach of Hoyer et al. [12]. For the remaining parts, we use the same set-up as for the MMA baselines - i.e. the cone filter with a radius of 2 pixels, and the material parameters set to $E_{0}=1.0$, $E_{min}=1e-9$ and $\nu=0.3$.
@@ 4 Results
We found that while optimizing topology without gradients, parameterizing the problem using the latent space of an autoencoder speeds up the process by at least one order of magnitude, closing the gap to gradient-based optimizers reported in the literature. Due to the choice of neural architecture, we demonstrate that the method shows significant robustness and generality, unlike elsewhere in the literature. These findings are demonstrated in Figure 2, which compares our LBAE and VAE architectures against the conventional baseline approach without reparameterization (pixel parameterization). In this example, the designs were optimized with BIPOP-CMA-ES (for complete results, including other optimizers, see the appendix). The performance is quantified as a fraction of problems (the cumulative probability) vs. the relative error of the final solution with respect to the MMA solution – considered here as the ground truth.
Figure 2: Cumulative probability (fraction of problems) vs relative error w.r.t. the MMA solution, for two variants of latent parameterization models - Variational Autoencoder (VAE, latent space dimensionality: 64), and Latent Bernoulli Autoencoder (LBAE, latent space dimensionality: 256), compared against conventional pixel parameterization. The designs were optimized with a BIPOP-CMA-ES optimizer. Note that the shown distributions do not add up to 1 due to the cut-off at 300%. Best, typical, and worst designs correspond to the best, median, and worst random initializations for each problem.
The results are grouped by problem set (in-distribution, and out-of-distribution problems). To investigate the robustness of our method, we also analyze the results considering only subsets of runs - for each design problem out of 15 randomly initialized optimization runs, we select the best, median, and worst run, to calculate the cumulative probability, as shown in 2 in columns 2-4 respectively. In all cases, the LBAE reparameterization significantly outperforms conventional black-box optimization within the prescribed evaluation budget (2000 FEM simulations), regardless of whether we choose all, best, median, or worst runs. In addition, the worst runs of LBAE remain significantly better than the best runs of the conventional pixel parameterization method, proving the overall robustness of this approach.
@@
Comparing LBAE architecture with the VAE, LBAE achieves significantly better performance overall, with the most significant difference visible on the subset of the worst runs, indicating on robustness of our approach. The VAE still achieves a significant margin over the conventional black-box optimization, thus proving the overall benefit of reparameterizing into the latent space, even with a more naive architecture. Nevertheless, the margins between LBAE and VAE indicate that the choice of architecture has a crucial influence on the performance.
Additionally, we find that the performance in terms of relative errors remains relatively intact whether it is evaluated on grayscale or thresholded designs. This is an important finding since historically many generative models struggled with producing representations with a significant presence of gray values, which after thresholding leads to disconnected structures (and thus a significantly deteriorated performance) [2]. However, in our case, the designs obtained with LBAE suffer from the gray-value problem to the same extent as the designs obtained with MMA.
Despite a significant advantage over the conventional pixel parameterization, the final objective values when compared to the gradient-based solutions can still be substantial – some designs show errors on the order of 100%, i.e. the compliance value is double the one found with MMA. For these problems, the imposed budget of 2000 FEM evaluations turns out insufficient to find a good solution. Therefore, unsurprisingly, the designs obtained without gradients remain inferior to those obtained with MMA.
@@
We investigate further the performance of our method by analyzing example designs.4. Figure 3 shows an example problem L-shaped bracket design problem from the out-of-distribution set. The example is representative of the most common cases in both problem sets – the relative errors achieved with LBAE with respect to the MMA solution are relatively low, with the worst solutions being within 10-30% worse than MMA solutions in terms of the relative error, and the best designs within 5-10%. The designs are visually quite different from the ground truth, and also different from each other - highlighting the stochastic nature of the gradient-free optimization. Note that for LBAE, the thresholding plays in favor of the model, leading to even lower relative error values as compared to the MMA design. In the case of VAE, however, the solutions are significantly oversimplified, and lacking detail, which results in significantly worse performance.
@@
Figure 4: An example of a challenging out-of-distribution problem: a ’staircase’ design with Z-shaped distributed load, with clam support a the bottom edge.
Figure4 shows an example of a more challenging out-of-distribution problem, with a Z-shaped distributed load, requiring a more intricate arrangement of the material. The variation in performance is relatively large, with the worst LBAE design yielding 132% higher compliance compared to the baseline. The designs generated with VAE achieve orders of magnitude worse performance. As in the previous problem (Figure 3), the designs obtained with VAE are distinctly oversimplified, indicating on limited capacity of the VAE architecture as compared to LBAE.
Interestingly, thresholding affects the design performance more drastically, and inconsistently as compared to the first example in Figure 3. The reason is the connectivity of the designs, and more specifically how it is affected by the thresholding step. Noteworthy, some of the designs appear to perform much better after thresholding e.g. typical design in LBAE outperforms the design of MMA. The reason for that is that the design obtained with MMA does suffer to some extent from intermediate ’gray-value’ densities present at the elements with prescribed external forces. As a result of thresholding, some of these externally loaded elements become void of material, altering the connectivity and increasing significantly the compliance value of the MMA design. In that case, the placement of the material within the elements subject to external loads in the LBAE design turns out more favorable, and robust to thresholding, leading to better performance.
Finally, all designs obtained with the conventional black-box optimization approach with pixel parameterization were so far from convergence, that we did not include them in the plots. This behavior was observed for all the tested problems in both problem sets, which indicates that the budget of 2000 FEM evaluations is simply insufficient for the conventional black-box optimizer to converge. This observation is in good agreement with the values reported in literature [8], stating typical numbers in order of 10,000 - 100,000 FEM iterations until convergence.
Heat Conduction problem
@@
We set up a simple heat-conduction example following that from [58]: there is a uniformly distributed heat source across the square design domain, and a heat sink in the middle of one of the edges, with the length of 0.2 of the edge length. The results obtained with our methods are compared to MMA in Figure 5. Even though our model has never been exposed to thermal problems of any kind, it can produce a solution, with a final objective (1.70) relatively close to that obtained with MMA (1.28). The baseline VAE achieves a better score than LBAE, however the design is less reminiscent of the one obtained with MMA (ground truth).
@@
Figure 5: Thermal compliance problem: comparison of latent optimization with CMAES and Latent Bernoulli Autoencoder (LBAE) vs MMA, unprocessed solution as well as thresholded
@@
Additionally, we consider objectives evaluated on thresholded designs, where the densities are rounded to 0 or 1, to eliminate the gray values, while preserving the volume fraction. Thresholding highlights the challenge of this design problem – the objective is very sensitive to the intermediate density values, making it an interesting benchmark. Due to the uniform heat source, the optimizers attempt to cover as much surface as possible, even at the cost of distributing intermediate densities of material, leading to gray values. As a result, the value of the objective evaluated on the thresholded design is two orders of magnitude larger, as compared to the objective evaluated on the design that has not been thresholded. Comparing the thresholded designs, latent optimization (LBAE with CMAES) leads to much worse performance, with the objective over four times larger as opposed to that obtained with thresholded MMA design. The thresholded design obtained with VAE yields much higher (worse) objective when compared to LBAE.
@@
This problem also offers a good opportunity to demonstrate some advantage of the gradient-free approach – instead of optimizing for the continuous density design (as required by the gradient-based approach), we can easily change the objective to optimize directly for the thresholded design objective. In other words, instead of optimizing for the objective evaluated with continuous density values, and thresholding in a postprocessing step, we can provide the optimizer with objective values evaluated directly on the thresholded designs. Optimizing for this objective leads to a much better loss value, as shown in 6, which drops almost fourfold compared to that in 5, and is now only 20% higher than the thresholded MMA design.
@@
Figure 6: Thermal compliance problem: latent optimization with CMAES and Latent Bernoulli Autoencoder (LBAE), optimized for thresholded objective, compared with MMA optimized for unthresholded objective and, thresholded afterwards, in the postprocessing step.

### v2

@@ 3.2 Benchmarking set-up
We develop and benchmark our method considering a standard topology optimization problem, namely compliance minimization with a volume constraint. This problem can be easily solved using gradient-based approaches, so it is not the ultimate application of the proposed method. Nevertheless, it provides a good assessment for the development of the method because, first, the “ground truth” solutions can be easily obtained using a gradient-based approach, allowing us to verify ’how far’ our method is from a good solution. Secondly, these problems are relatively cheap to evaluate, as opposed to design problems based on more complex physics, such as the fracture problem presented towards the end of the article.
@@
To provide statistically meaningful performance measures, we test our optimization framework on two sets of optimization problems. The first set contains 25 in-distribution design problems (sets of boundary conditions). These problems were generated in the same way as the problems seen during the training (i.e., with random point loads and point supports). The second set contains 25 out-of-distribution problems, generated using the approach of Hoyer et al. [19] (the same as the testing dataset). These problems include distributed loads and supports, therefore they provide a better insight into the generalization of the model.
@@
To quantify the performance of different model and optimizer configurations across the whole problem set, for each problem, we normalize the compliance value with the compliance of the MMA solution (considered as the ’ground truth’), and in this way calculate the relative error. Furthermore, to account for stochastic factors within the optimization process (e.g., initialization of the optimizer), the optimization process for each configuration and each problem is repeated with 15 different random seeds. Performance is evaluated for grayscale and binary (thresholded) designs. For thresholding we apply the simplest strategy from [66] in a post-processing step. In that case, the relative error is calculated with respect to the value obtained for the thresholded MMA design.
@@
For gradient-free optimizer, we use Covariance Matrix Adaptation Evolutionary Strategies (CMA-ES) [60, 67] and its Bi-Population variant (BIPOP-CMA-ES) [68]. The CMA-ES and its variants (such as BIPOP-CMA-ES) are currently regarded as state-of-the-art gradient-free optimization algorithms [14]. Furthermore, CMA-ES was considered in a similar setting [69]. Notwithstanding, we considered several alternative gradient-free optimizers to ensure that we selected the most favorable one for our problems, namely Particle Swarm Optimization [70], Differential Evolution [71], Evolutionary Strategies meta-optimizer [72], and Simple Genetic Algorithm [73]. However, even with hyperparameter tuning, these optimizers exhibited considerably worse performance than CMA-ES within the assigned budget.
@@
We use the density-based formulation of topology optimization, which is the most widely adopted approach, although numerous other formulations are available [1]. The volume constraint is enforced following the approach of Hoyer et al. [19], by mapping the outputs of the neural network $\tilde{x}$ to density values $x$ with the following sigmoid transformation:
$x=\frac{1}{1+\exp{\left(\tilde{x}-b(\tilde{x},V_{0})\right)}}$ $s.t:\int_{\Omega}\rho=V_{0}$ (2)
where the constant $b(x,V_{0})$ is found using a bisection algorithm, such that the final volume of the density field over the design domain $\Omega$ satisfies the prescribed volume $V_{0}$.
For the remaining parts, we use the same set-up as for the MMA baselines - i.e. the cone filter with a radius of 2 pixels, and the material parameters set to $E_{0}=1.0$, $E_{min}=1e-9$ and $\nu=0.3$.
@@ 4 Results
We found that while optimizing topology without gradients, parameterizing the problem using the latent space of an autoencoder speeds up the process by at least one order of magnitude, closing the gap to gradient-based optimizers reported in the literature. Due to the choice of neural architecture, we demonstrate that the method shows significant robustness and generality, unlike elsewhere in the literature. These findings are demonstrated in Figures 2 and 3, which compare our LBAE and VAE architectures against the conventional baseline approach without reparameterization (pixel parameterization). In this example, the designs were optimized with BIPOP-CMA-ES (for complete results, including other optimizers, see the appendix). Performance is quantified as a fraction of problems (the cumulative probability) vs. the relative error of the final solution with respect to the MMA solution – considered here as the ground truth.
The results are grouped by problem set (in-distribution, and out-of-distribution problems). To investigate the robustness of our method, we also analyze the results considering only subsets of runs – for each design problem out of 15 randomly initialized optimization runs, we select the best, median, and worst run, to calculate the cumulative probability, as shown in Figures 2 and 3 in columns 2-4, respectively. In all cases, the LBAE reparameterization significantly outperforms conventional black-box optimization within the prescribed evaluation budget (2000 FEM simulations), regardless of whether we choose all, best, median, or worst runs. In addition, the worst runs of LBAE remain significantly better than the best runs of the conventional pixel parameterization method, proving the overall robustness of this approach.
Figure 2: Performance of unprocessed designs: cumulative probability (fraction of problems) vs relative error w.r.t. the MMA solution, for two variants of latent parameterization models – Variational Autoencoder (VAE, latent space dimensionality: 64), and Latent Bernoulli Autoencoder (LBAE, latent space dimensionality: 256) – compared against conventional pixel parameterization. The designs were optimized with a BIPOP-CMA-ES optimizer. Note that the shown distributions do not add up to 1 due to the cut-off at 300%. Best, typical, and worst designs correspond to the best, median, and worst random initializations for each problem.
@@
When comparing LBAE architecture with the VAE, LBAE achieves significantly better performance overall, with the most significant difference visible on the subset of the worst runs, indicating the robustness of our approach. The VAE still achieves a significant margin over the conventional black-box optimization, thus proving the overall benefit of reparameterizing into the latent space, even with a more naive architecture. Nevertheless, the margins between LBAE and VAE indicate that the choice of architecture has a crucial influence on the performance.
Figure 3: Performance of thresholded designs: cumulative probability (fraction of problems) vs relative error w.r.t. the MMA solution, calculated for thresholded designs for two variants of latent parameterization models – Variational Autoencoder (VAE, latent space dimensionality: 64), and Latent Bernoulli Autoencoder (LBAE, latent space dimensionality: 256) – compared against conventional pixel parameterization. The designs were optimized with a BIPOP-CMA-ES optimizer. Note that the shown distributions do not add up to 1 due to the cut-off at 300%. Best, typical, and worst designs correspond to the best, median, and worst random initializations for each problem.
Additionally, we find that the performance in terms of relative errors remains relatively intact whether it is evaluated on grayscale or thresholded designs. This is an important finding, as historically many generative models struggled to produce representations with a significant presence of gray values, which after thresholding leads to disconnected structures (and thus a significantly deteriorated performance) [6]. However, in our case, the designs obtained with LBAE suffer from the gray-value problem to the same extent as the designs obtained with MMA.
Despite a significant advantage over the conventional pixel parameterization, the final objective values when compared to the gradient-based solutions can still be substantial – some designs show errors on the order of 100%, i.e. the compliance value is double the one found with MMA. For these problems, the imposed budget of 2000 FEM evaluations turns out to be insufficient to find a good solution. Therefore, unsurprisingly, the designs obtained without gradients remain inferior to those obtained with MMA.
@@
We investigate further the performance of our method by analyzing example designs. Figure 4 shows an example problem of an L-shaped bracket design from the out-of-distribution set. The example is representative of the most common cases in both problem sets – the relative errors achieved with LBAE with respect to the MMA solution are relatively low, with the worst solutions being within 10-30% worse than MMA solutions in terms of the relative error, and the best designs within 5-10%. The designs are visually quite different from the ground truth and also different from each other – highlighting the stochastic nature of the gradient-free optimization. Note that for LBAE, the thresholding plays in favor of the model, leading to even lower relative error values compared to the MMA design. In the case of VAE, however, the solutions are significantly oversimplified and lacking detail, which results in significantly worse performance.
@@
Figure 5 shows an example of a more challenging out-of-distribution problem, with a Z-shaped distributed load, requiring a more intricate arrangement of the material. The performance variation is relatively large, with the worst LBAE design yielding 132% higher compliance compared to the baseline. The designs generated with VAE achieve orders of magnitude worse performance. As in the previous problem (Figure 4), the designs obtained with VAE are distinctly oversimplified, indicating the limited capacity of the VAE architecture as compared to LBAE.
Interestingly, thresholding affects design performance more drastically and inconsistently compared to the first example in Figure 4. The reason is the connectivity of the designs and, more specifically, how it is affected by the thresholding step. Noteworthy, some of the designs appear to perform much better after thresholding e.g. typical design in LBAE outperforms the design of MMA. The reason for this is that the design obtained with MMA does suffer to some extent from intermediate ’gray-value’ densities present at the elements with prescribed external forces. As a result of thresholding, some of these externally loaded elements become void of material, altering the connectivity and significantly increasing the compliance value of the MMA design. In that case, the placement of the material within the elements subject to external loads in the LBAE design turns out to be more favorable and robust to thresholding, leading to better performance.
Figure 5: An example of a challenging out-of-distribution problem: a ’staircase’ design with Z-shaped distributed load, with clamped support a the bottom edge.
Finally, all designs obtained with the conventional black-box optimization approach with pixel parameterization were so far from convergence that we did not include them in the plots. This behavior was observed for all the tested problems in both problem sets, which indicates that the budget of 2000 FEM evaluations is simply insufficient for the conventional black-box optimizer to converge. This observation is in good agreement with the values reported in the literature [13], stating typical numbers in order of 10,000 - 100,000 FEM iterations until convergence.
Model limitations: heat conduction problem
@@
We set up a simple heat-conduction example following the setup of [74]: there is a uniformly distributed heat source across the square design domain, and a heat sink in the middle of one of the edges, with the length of 0.2 of the edge length. The results obtained with our methods are compared to MMA in Figure 6. Although our model has never been exposed to thermal problems of any kind, it can produce a solution, with a final objective (1.70) relatively close to that obtained with MMA (1.28). The baseline VAE achieves a significantly worse score (28.65) with a design that lacks finer features, indicating possible limitations regarding VAE’s expressivity.
@@
Figure 6: Thermal compliance problem: comparison of latent optimization with CMAES and Latent Bernoulli Autoencoder (LBAE) vs MMA, unprocessed and thresholded solutions. The thresholded objective (bottom plot) was evaluated by thresholding the designs at each iteration, and evaluating the objective. The optimization, however, was carried out using unthresholded designs (top row).
@@
Additionally, we consider objectives evaluated on thresholded designs, where the densities are rounded to 0 or 1, to eliminate the gray values, while preserving the volume fraction. Thresholding highlights the challenge of this design problem – the objective is very sensitive to the intermediate density values, making it an interesting benchmark. Due to the uniform heat source, the optimizers attempt to cover as much surface as possible, even at the cost of distributing intermediate densities of the material, leading to gray values. As a result, the value of the objective evaluated on the thresholded design is two orders of magnitude larger, as compared to the objective evaluated on the design that has not been thresholded. When comparing the thresholded designs, latent optimization (LBAE with CMAES) leads to much worse performance, with the objective over four times larger than that obtained with the thresholded MMA design. The thresholded design obtained with VAE yields a much higher (worse) objective than LBAE.
@@
This problem also offers a good opportunity to demonstrate some advantages of the gradient-free approach – instead of optimizing for the continuous density design (as required by the gradient-based approach), we can easily change the objective to optimize directly for the thresholded design objective. In other words, instead of optimizing for the objective evaluated with continuous density values, and thresholding in a postprocessing step, we can provide the optimizer with objective values evaluated directly on the thresholded designs. Optimizing for this objective leads to a much better loss value, as shown in Figure 7, which drops almost four times compared to that of Figure 6, and is now only 20% higher than the thresholded MMA design.
@@
Figure 7: Thermal compliance problem: latent optimization with CMA-ES and Latent Bernoulli Autoencoder (LBAE), optimized for thresholded objective, compared with MMA optimized for un-thresholded objective and, thresholded afterward, in the postprocessing step.
Validation: optimizing structure undergoing brittle fracture
Our main hypothesis for this work is that improving gradient-free topology optimization via neural reparameterization can lead to better results for problems involving non-linear path-dependent physics, with non-smooth objective landscapes, such as brittle fracture. Although topology optimization with consideration of fracture has been addressed in several works [75, 10, 9, 15, 35, 76], it remains one of the most challenging problems in structural optimization. The main source of difficulty is the nonlinearity and path-dependence inherent to fracture, which make the objective function highly sensitive to design changes [9, 15]. This poses significant difficulties for gradient-based methods and opens up opportunities for a gradient-free approach. To our knowledge, this is the first time that gradient-free optimization has been applied to optimize structures undergoing brittle fracture. We conjecture that the curse of dimensionality together with the high computational cost of finite element analysis in each iteration has made it unfeasible to use gradient-free approaches in these topology optimization problems. Yet, the proposed gradient-free neural topology optimization strategy might make it feasible, if not even better than gradient-based topology optimization strategies for these problems due to the non-smoothness of the objective landscape.
For modeling fracture, we use a phase-field formulation [77, 78], as already adopted in the context of topology optimization [75]. Compared with alternative approaches such as X-FEM, phase-field facilitates the handling of initiation, propagation, and merging of cracks. It also allows for solving on a fixed mesh which makes it relatively easy to implement and integrate with topology optimization frameworks.
The main principle behind the phase-field method is to model the crack in diffusive form by using an auxiliary scalar field $d\in[0,1]$, such that the crack surface can be approximated using the following elliptic functional [79, 77]:
[…]
where $\gamma$ is the crack surface density function, and can be expressed in terms of the phase-field variable $d$ as:
$\gamma(d,\nabla d)=\frac{1}{2l}d^{2}+\frac{l}{2}\|\nabla d\|^{2}$ (4)
The parameter $l$ is the characteristic length scale that controls the regularization of the problem (crack smearing). With such an approximation, the energy functional over phase-field $d$ and displacement field $\mathbf{u}$ can be formulated as follows:
[…]
where $G_{c}$ is the Griffith-type energy release rate for a given material, and $\psi$ is the isotropic elastic energy density. The elastic energy density term undergoes an additive decomposition into compressive and tensile components, as shown in Equation (6).
$\psi(\mathbf{\epsilon}(\mathbf{u}),d)=\psi^{-}(\epsilon(\mathbf{u}))+g(d)\psi^{+}(\mathbf{\epsilon}(\mathbf{u}))$ (6)
where the compressive and tensile parts are defined using spectral decomposition of the strain tensor [77]:
$\psi^{\pm}=\frac{\lambda}{2}\langle tr[\epsilon(\mathbf{u})]\rangle^{2}_{\pm}+\mu tr\left[\sum_{i=1}^{dim}\langle\epsilon^{i}(\mathbf{u})\rangle_{\pm}\mathbf{n^{i}}\otimes\mathbf{n^{i}}\right]^{2}$ (7)
Here $\langle\rangle_{\pm}$ denote the Macaulay brackets, defined as $\langle x\rangle_{+}=\max(0,x)$ and $\langle x\rangle_{-}=\min(0,x)$. The terms $\mathbf{n^{i}}$ and $\epsilon^{i}(\mathbf{u})$ are respectively the eigenvectors and eigenvalues of the strain tensor, and $g(d)$ is the degradation function, which takes the form: $g(d)=(1-d)^{2}$. The purpose of spectral decomposition into tensile and compressive parts is to ensure that the crack growth can be induced only in tension – as can be seen in Equation (7), the degradation function term affects only the tensile term.
The strong form can be obtained by minimizing the energy stated in Equation (5), such that $\delta\Pi=0$. This results in the following coupled PDE system:
$\nabla\cdot\mathbf{\sigma}(\mathbf{u},d)=\mathbf{0}$ in $\text{in }\Omega$ (8)
$\frac{G_{c}}{l}\left(d-l^{2}\Delta d\right)=2(1-d)\mathcal{H}$ in $\text{in }\Omega$ (9)
$\mathbf{n}\cdot\mathbf{\sigma}(\mathbf{u},d)=\mathbf{t}$ on $\text{on }\partial\Omega^{t}$ (10)
$\mathbf{u}=\mathbf{u^{D}}$ on $\text{on }\partial\Omega^{u}$ (11)
$\mathbf{n}\cdot\nabla d=0$ on $\text{on }\partial\Omega$ (12)
Equation (8) represents the quasi-static force balance, while Equation (9) governs the phase field evolution. Here $\mathcal{H}$ is the history variable, defined as $\mathcal{H}(\mathbf{x},t)=\max_{\tau\in[0,t]}\left[\psi_{e}^{+}(\mathbf{x},t)\right]$, which enforces irreversibility of the damage or, in other words, prevents crack healing. In Equation (11) the term $\mathbf{n}$ corresponds to the surface normal vector, $\mathbf{t}$ represents the traction applied on the boundary $\partial\Omega^{t}$. The term $\mathbf{u}^{D}$ corresponds to the Dirichlet boundary conditions prescribed on the boundary $\partial\Omega^{u}$.
This strong form gives rise to a weak form that is discretized using the finite element method (for details, see e.g. [78, 15, 9]) and solved numerically. Due to the path dependence, the loading is applied in increments. The solutions for $\mathbf{u}$ and $d$ are found by solving the coupled systems of the FEM equations using the standard iterative staggered scheme, which is the most common approach [15, 10, 9]. Furthermore, the displacement equations, which are non-linear due to material degradation, are solved using the iterative Newton method. More details on the parameters used in this fracture simulation can be found in the appendix.
The optimization problem is formulated as maximizing the total external work, with maximum volume constraint, as shown in Equation (14). The formulation is akin to that posed in [9, 10].
[…]
The parameterization of material properties is based on the SIMP method, following the approach of [15]. While using SIMP comes with certain limitations [15], such as non-physical ‘gray-values’, it remains a common practice in the topology optimization community [76, 15]. To regularize the solution and prevent checkerboard patterns, we apply a standard cone filter [18]. The boundary conditions for our benchmark problem correspond to a simple cantilever beam, as shown in Figure 8, that is based on examples of [9] and [15].
Figure 8: Illustration of the boundary conditions of the cantilever beam problem. The left edge is clamped, while at the right edge, the load acting downwards is applied in terms of the prescribed displacement $u^{D}$. The displacement is applied in several increments until it reaches a maximum of 0.5 mm.
We solve this optimization problem using our proposed approach of gradient-free neural topology optimization and compare it with a traditional solution method, i.e. using a gradient-based MMA algorithm. Importantly, we use our LBAE model without retraining, i.e. it has been trained only once on the dataset comprised of designs optimized for compliance. Furthermore, to match the shape and size of the discretized domain of our benchmark (64x128 elements) we resize the output of the LBAE model using standard bi-linear interpolation. We set the CMA-ES population size to 16, which means that in each iteration 16 candidate solutions are evaluated (in parallel). For the baseline gradient-based solution we use sensitivities obtained with automatic differentiation based on JAX [80, 81].
Both gradient-based and gradient-free approaches were evaluated using two different initialization strategies. In the first scenario, we use the default initialization, i.e. for MMA the initial design was set to a uniform density field satisfying the volume fraction constraint, while the initial CMA-ES distribution was initialized using the default $\mathbf{0}$ mean. In the second strategy, the optimizers were initialized using a design optimized for minimum compliance using a linear elasticity model. In the case of CMA-ES, this initial design was encoded into the latent vector $\mathbf{z}$ using the encoder, and the latent vector was used to set the mean of the initial distribution. Additionally, for the gradient-free optimizer, we tested different initial values for $\sigma_{0}$, which controls the spread in the initial distribution.
The results are summarized in Figure 9, which presents the final design and the corresponding objective values. As shown in the figure, gradient-free optimization outperforms the gradient-based approach by a significant margin (around 30%) in all tested configurations. Investigating the designs obtained with CMA-ES, it can be seen that in all cases the optimizer managed to identify a better design strategy, reinforcing the upper left corner, where the tensile loads are maximum, and in this way postponing fracture. Following this strategy, the designs are characterized by an asymmetric layout – a typical feature observed in designs optimized with consideration of fracture [9, 15].
Figure 9: Comparison of final designs obtained with our proposed gradient-free approach - gradient-free CMA-ES with LBAE reparameterization compared against the baseline designs obtained with gradient-based MMA optimizer. The results were obtained for two different initialization strategies: default (uniform for MMA and $\mathbf{0}$ mean for CMA-ES) as well as starting from the design optimized for compliance governed by linear elasticity (in that case, the mean of CMA-ES distribution is initialized from the compliance design encoded into a latent vector). Furthermore, CMA-ES was tested for 3 different settings of the initial variance parameter $\sigma_{0}$ for the initial population distribution.
On the contrary, the gradient-based optimizer struggled to identify this strategy, due to getting stuck in local optima. In the case of initialization from the compliance design (Figure 9e), the optimizer gets stuck in a local minimum which exploits the SIMP parameterization. In an attempt to reduce the damage due to fracture, the optimizer lowers the density values throughout the whole domain, leading to lower stiffness and lower stresses, in such a way attempting to prevent damage. This leads to the "gray values" design with a relatively low stiffness (and thus lower external work needed to reach the prescribed displacement), which results in mediocre overall performance. However, while attempting to increase the density values, damage is induced in the design, resulting in a ‘barrier’ in the loss landscape, which the optimizer is unable to cross.
In the second case (Figure 9a) the evidence of the optimizer getting stuck in the local minimum is more subtle, yet it can be found by investigating the force-displacement curves shown in Figure 10. Unlike the design obtained with CMA-ES (Figure 10d), which is reinforced in the upper left corner, the structure obtained with MMA (Figure 10a) is allowed to fracture in the upper corner relatively quickly, and instead, the optimizer attempts to reinforce the remaining intact part (lower left corner) - or, in other words, improve the post-fracture response of the design. This manifests itself in the force-displacement curve as a drop in the reaction force at the displacement of $u^{D}=0.3$ mm. However, this approach leads to a suboptimal design with around 30% lower strain energy absorption, as opposed to the CMA-ES solution. As can be seen from the history of the force-displacement curves, this characteristic saw-tooth response pattern is present starting in the early iterations. It appears that the MMA optimizer is not able to change the response of the design, and gets stuck in a suboptimal reinforcement strategy.
Figure 10: Comparison of optimization with fracture using gradient-based MMA vs gradient-free CMA-ES with LBAE reparameterization. Pictures a) and d) represent the best designs with the phase-field damage marked in red. For CMA-ES in each iteration, the whole population is evaluated, i.e. 16 designs are simulated in parallel.